A steamer goes downstream in a river, covers the distance between 2 towns in 15
hours. Coming back upstream, it covers this distance in 20 hours. The speed of the water is 3 km/h,find the distance between two towns.

Answer by lwsshak3(11628)

A steamer goes downstream in a river, covers the distance between 2 towns in 15
hours. Coming back upstream, it covers this distance in 20 hours. The speed of the water is 3 km/h,find the distance between two towns.
**
let x=speed of steamer in still water
x+3=speed of steamer downstream
x-3=speed of steamer upstream
distance=travel time*speed (same for both directions)
..
15(x+3)=20(x-3)
15x+45=20x-60
5x=105
x=21
distance=15(x+3)=15*24=360
distance between two towns=360 km